OpenStudy (anonymous):

Last problem! Help for medal. I need help with just one more inequality equation... x^2 - x - 15 > 15

OpenStudy (anonymous):

x<-5 or x>6

OpenStudy (anonymous):

Would you like me to explain?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

sure

OpenStudy (usukidoll):

@_greatmath7 thanks for providing an explanation in addition to the direct answer. :)

OpenStudy (anonymous):

Let's solve your inequality step-by-step. x^2-x-15>15 Let's find the critical points of the inequality. (Subtract 15 from both sides): x^2-x-15-15=15-15 You then receive: x^2-x-30=0 (Factor left side of equation):(x+5)(x-6)=0 (Set factors equal to 0):x+5=0 and x-=0

OpenStudy (anonymous):

thnx @UsukiDoll

OpenStudy (anonymous):

x-6=0

OpenStudy (anonymous):

Oh I see. Thank you very much! :)

OpenStudy (usukidoll):

o_0 \[\LARGE x^2 - x - 15 > 15\] first we subtract 15 from both sides \[\LARGE x^2 - x - 15-15 > 15-15\] \[\LARGE x^2 - x - 30 > 0\] then we have to factor. our last term is -30 and the middle term is -1 6 x -5 is the only combination that can work for this quadratic equation because 6 x -5=-30 6-5 = 1 Therefore \[\LARGE (x-6)(x+5) > 0\] now we solve \[\LARGE (x-6)(x+5) > 0\] by splitting this up into two cases \[\LARGE (x-6) > 0\] \[\LARGE (x+5) > 0\] solving for an inequality is the same as solving for an equation \[\LARGE x-6 > 0\] \[\LARGE x-6+6 > 0+6\] \[\LARGE x > 6\] \[\LARGE x+5 > 0\] \[\LARGE x+5-5 > 0-5\] \[\LARGE x > -5\] time to switch the inequality sign \[\LARGE x < -5\]

OpenStudy (usukidoll):

on the number line since we have < we have an open circle